### v2

# Developmental Symmetry-Loss: A Free-Energy Perspective on Brain-Inspired Invariance Learning

Abstract
We propose Symmetry-Loss, a brain-inspired algorithmic principle that enforces invariance and equivariance through a differentiable constraint derived from environmental symmetries. The framework models learning as the iterative refinement of an effective symmetry group $\hat{G}$, paralleling developmental processes in which cortical representations align with the world’s structure. By minimizing structural surprise, i.e. deviations from symmetry consistency, Symmetry-Loss operationalizes a Free-Energy–like objective for representation learning. This formulation bridges predictive-coding and group-theoretic perspectives, showing how efficient, stable, and compositional representations can emerge from symmetry-based self-organization. The result is a general computational mechanism linking developmental learning in the brain with principled representation learning in artificial systems.
Motivation
Brains do not learn by passively fitting data; they develop internal models that reflect the symmetries and regularities of the world. From early visual processing to higher cognition, neuronal populations gradually acquire invariances—such as translation, reflection, and compositional structure—that stabilize perception and support generalization (DiCarlo2012; Bertamini2018; Fumarola2022). This developmental process can be interpreted as a progressive alignment of internal symmetries with environmental dynamics (Higgins2022; Cooray2024).
Recent theoretical frameworks, including the Free-Energy Principle and predictive coding, describe this alignment as the minimization of surprise or prediction error (Friston2010; Parr2019). Similarly, symmetry-based representation learning in machine intelligence (Higgins2022) and cognitive-developmental theories (Lake2017; Tenenbaum2011) both emphasize structure discovery through transformation consistency.
In this work, we propose Developmental Symmetry-Loss, a differentiable mechanism that operationalizes this principle within modern learning systems. The method penalizes deviations from group-consistent transformations and iteratively refines an effective symmetry group $\hat{G}$, which stabilizes representational dynamics. Conceptually, this procedure mirrors how cortical representations specialize through symmetry preservation and symmetry breaking during development (Fumarola2022; Keck2025).
By interpreting symmetry consistency as a form of structural free-energy minimization, our approach links predictive-coding theories in neuroscience with invariant learning in machine intelligence (Friston2010; Whittington2019). The resulting framework provides both (i) a theoretical bridge between biological and artificial learning and (ii) a practical mechanism for stable, energy-efficient, and interpretable representation formation.
In summary, Developmental Symmetry-Loss proposes that learning invariances through iterative symmetry refinement is not merely an engineering strategy—but an algorithmic reflection of how the brain itself learns to model its world.
Symmetry Refinement: Theory and Formulation
Learning in biological and artificial systems can be viewed as the process of aligning internal transformations with the symmetries of the environment. We formalize this idea through a differentiable objective that penalizes representational deviations from group-consistent transformations.
Prerequisites
Foundations: Actions, invariants.
Let $G$ denote a candidate topological group, and $\mathcal{F}$ a topological $G$-modul. We denote by $\mathcal{F}^{G}:=\{f\in\mathcal{F}\,:\,g.f=f~~~\forall g\in G\}$ the zeroth cohomology group of $G$ with coefficients in $\mathcal{F}$. We call its elements symmetries. In the present work, we are only interested in the case where $\mathcal{F}$ is some class of functionals of a latent space $\mathcal{X}$, i.e. $\mathcal{F}=\mathcal{F}(\mathcal{X})$.
We use the basic algebraic geometry setting (shafarevich1994basic). Although algebraic geometry is not essential for understanding the framework, it provides a solid and rigorous foundation that ensures the formulation remains technically accurate. Let the latent space $\mathcal{X}\subseteq\mathbf{A}^{n}:=\mathbf{K}^{n}$ be an affine variety over $\mathbf{K}\in\{\mathbf{R},\mathbf{C}\}$, and let $G$ be an algebraic group that acts polynomial on $\mathcal{X}$. This means $\mathcal{X}$ is a $G$-variety. In this context, we are interested in the class of polynomial functionals also known as regular functionals. We denote these functionals as $\mathcal{F}=\mathbf{K}[\mathcal{X}]=\{f:\mathcal{X}\to\mathbf{K}\,:\,f\;\mathrm{is\;regular}\}$.
The action of $G$ on $\mathcal{X}$ makes $\mathcal{F}$ to a topological $G$-module, defined by $(g\cdot f)(x):=f(g^{-1}x)$, where $g\in G$, $f\in\mathcal{F}$, and $x\in\mathcal{X}$ (kraft1985geometrische). We are particularly interested in finding symmetries in $\mathcal{F}$, i.e. $\mathcal{F}^{G}\subseteq\mathcal{F}$. In other words, we seek polynomial functionals of the latent space that remain invariant under the action of $G$. The ring structure of $\mathcal{F}^{G}$ is inherited by $\mathcal{F}$, also called the invariant ring.
Geometric invariant theory quotient.
From this point forward, let $G$ be a linear reductive group. When $G$ acts on a latent space $\mathcal{X}$, the first step is to examine the quotient map $\pi:\mathcal{X}\to\mathcal{X}//G$ induced by the action. Understanding this map is essential, as it reveals the structural properties of the action itself.
In the polynomial setting, the quotient map takes the form of the geometric invariant theory quotient (GIT quotient):
$\pi:\mathcal{X}\to\mathcal{X}//G,\;x\mapsto\left(\chi_{1}(x),\ldots,\chi_{m}(x)\right),$
where $\mathcal{X}//G:=\mathrm{Spec}\left(\mathcal{F}^{G}\right)\subseteq\mathbf{A}^{m}$.
This is a categorical quotient and the functionals $\chi_{1},\ldots,\chi_{m}\in\mathcal{F}^{G}$ form a complete fundamental system of symmetries, meaning that $\mathbf{K}[\chi_{i}:i=1,\ldots,m]=\mathcal{F}^{G}$. Notably, the function $\pi$ serves to identify orbit closures; that is, for any points $v,x\in\mathcal{X}$, we have $\overline{Gx}=\overline{Gv}$ if and only if $\pi(x)=\pi(v)$ (kraft1985geometrische; mumford1994geometric).
This elegantly illustrates how symmetries of $\mathcal{F}$ are linked to the group’s action on the domain space.
Target-free symmetry refinement and disentanglement
Symmetry-Loss functional.
Let $\Omega$ denote the input/sample space, $\mathcal{X}$ a latent space and $\phi_{\theta}:\Omega\rightarrow\mathcal{X}$ be a parameterized representation map.
Let $G$ denote a candidate group of transformations acting on the latent space $\mathcal{X}$, and let $\chi_{1},\ldots,\chi_{m}\in\mathcal{F}(\mathcal{X})$ be its complete fundamental system of symmetries.
For $v_{0}\in\Omega$ let
$\left(\chi_{1}(\phi_{\theta}(v_{0})),\ldots,\chi_{m}(\phi_{\theta}(v_{0}))\right)=:\left(y_{1},\ldots,y_{m}\right)\in\mathbf{A}^{m}.$
Since the $\chi_{i}$’s identify orbits (closures), deviations from this relation quantify a violation of orbit-alignment.
We therefore define the Symmetry-Loss relative $v_{0}$ as
$\mathcal{L}_{\mathrm{sym}}(\phi_{\theta};G,v_{0})=\sum_{i=1}^{m}\,\mathbf{E}_{\omega\sim\mathcal{D}}\Big[\|\,y_{i}-\chi_{i}\left(\phi_{\theta}(\omega)\right)\|^{2}_{2}\Bigg],$
where $\mathcal{D}$ is the empirical data distribution over $\Omega$, and $G$ the current hypothesis of environmental symmetries. If the context is clear, we will drop the $v_{0}$ in the notation.
Minimizing $\mathcal{L}_{\mathrm{sym}}$ (with respect to the model parameters $\theta$) enforces local consistency relative $v_{0}$ of representations under the group action, analogous to minimizing prediction error in predictive-coding theories. In settings lacking a fixed anchor element $v_{0}$, the symmetry–loss is minimized jointly over the model parameters and the latent variables $y_{i}$, allowing the system to self–organize reference points and thereby eliminate dependence on an externally chosen anchor.
We further define the orbit-loss of $x\in\mathcal{X}$ relative $v_{0}$ as
$\mathcal{L}_{\mathcal{O},v_{0}}(\phi_{\theta};G,x)=\sum_{i=1}^{m}\,\|\,y_{i}-\chi_{i}\left(x\right)\|^{2}_{2}.$
Structural surprise.
We interpret $\mathcal{L}_{\mathrm{sym}}$ as a measure of structural surprise: the mismatch between the symmetry predicted by the model relative $v_{0}$ and the symmetry expressed by the data distribution $\mathcal{D}$. Whereas classical free-energy minimization reduces sensory surprise, symmetry-loss reduces structural surprise by aligning internal transformations with external regularities.
The preceding discussion used algebraic–geometric language (varieties, invariant rings, GIT quotients) to clarify how group actions induce symmetry coordinates. In what follows we work in the smooth category: all spaces are smooth manifolds and maps are diffeomorphisms (or locally diffeomorphic in practice). Conceptually, the two regimes are parallel:
  • (i)
    invariant generators $(\chi_{i})$ provide orbit–separating coordinates in the algebraic case, while in the smooth case one may use separating invariants, sufficient statistics, or learned surrogates in the symmetry–loss;
  • (ii)
    categorical quotients $\mathcal{X}\!//G$ correspond to orbit spaces modulo $G$ in the smooth setting (up to regularity);
  • (iii)
    group actions, intertwiners, and orbit–alignment have identical formal roles.
The GIT intuition (quotients via invariants) is retained as a guiding picture, but all optimization and refinement steps are carried out in the smooth category.
Iterative Disentanglement by Symmetry Refinement.
In the absence of explicit targets, learning reduces to organizing representations so that they respect the intrinsic symmetries of the data. Rather than aligning outputs to fixed labels, the model aligns internal transformations to the structural regularities of the input space itself. This defines a target–free or self–supervised setting, in which the learning signal arises purely from deviations from orbit consistency.
Consider a representational space $\Omega\subset\mathbf{A}^{n}$ whose samples are subject to latent transformations. We assume that the observed variability in $\Omega$ arises from the (possibly partial or unknown) action of a group $G$. Our aim is to iteratively uncover a coordinate system in which this action becomes increasingly disentangled.
Let $\Gamma_{0}=\Omega\times\Omega$ denote the initial relation between inputs and their current representations. Given an input group $G_{0}$ acting on $\Omega_{1}\times\Omega$, we learn a diffeomorphism $\phi_{0}:\Omega\to\Omega_{1}$ such that the transformed graph
$\Gamma_{0}^{\prime}=\{\,(\phi_{0}(x),\,x)\mid x\in\Omega\,\}\subseteq\Omega_{1}\times\Omega$
approximates a single orbit of $G_{0}$, i.e.
$\Gamma_{0}^{\prime}\approx G_{0}\cdot(\phi_{0}(x_{0}),\,x_{0}),$
by minimizing $\mathcal{L}_{\mathrm{sym}}(\phi_{0};G_{0})$ for orbit consistency. After minimizing, a residual misalignment typically remains, indicating unmodelled latent symmetries. We therefore introduce a new latent space $\Omega_{2}$, a group $G_{1}$ acting on $\Omega_{2}\times\Omega$, and a further coordinate refinement $\phi_{1}:\Omega_{1}\to\Omega_{2}$, again minimizing $\mathcal{L}_{\mathrm{sym}}(\phi_{1};G_{1})$, such that the transformed graph
$\Gamma_{1}^{\prime}=\{\,(\phi_{1}(\phi_{0}(x)),\,x)\mid x\in\Omega\,\}\subseteq\Omega_{2}\times\Omega$
approaches a single orbit of $G_{1}$. Repeating this yields a hierarchy
$(\phi_{k},G_{k})_{k\geq 0},\qquad\Omega_{k+1}=\phi_{k}(\Omega_{k}),$
which defines an iterative disentanglement process. Each step refines the representation by reparametrizing coordinates so that the new latent manifold becomes increasingly symmetric.
Intertwiners and effective symmetries.
Each refinement step defines an equivariant morphism
$(\phi_{k}\times\mathrm{id}):(\Omega_{k}\times\Omega,\,G_{k}^{(k)})\longrightarrow(\Omega_{k+1}\times\Omega,\,G_{k}),$
where $G_{k}^{(k)}$ denotes the induced action of $G_{k}$ on $\Omega_{k}\times\Omega$ obtained by conjugation through $\phi_{k}\times\mathrm{id}$:
$G_{k}^{(k)}:=(\phi_{k}\times\mathrm{id})^{-1}\,G_{k}\,(\phi_{k}\times\mathrm{id})\subset\mathrm{Diff}(\Omega_{k}\times\Omega).$
This definition ensures that each map $\phi_{k}\times\mathrm{id}$ is an intertwiner between the group actions $(G_{k}^{(k)},\,\Omega_{k}\times\Omega)$ and $(G_{k},\,\Omega_{k+1}\times\Omega)$, satisfying
$(\phi_{k}\times\mathrm{id})\circ g^{(k)}=g\circ(\phi_{k}\times\mathrm{id}),\qquad g\in G_{k}.$
For any earlier index $j<k$, the corresponding action of $G_{j}$ can be transported to the level $\Omega_{k}\times\Omega$ via repeated conjugation,
$G_{j}^{(k)}:=(\phi_{k-1}\times\mathrm{id})\cdots(\phi_{j+1}\times\mathrm{id})\,G_{j}\,(\phi_{j+1}\times\mathrm{id})^{-1}\cdots(\phi_{k-1}\times\mathrm{id})^{-1},$
so that all induced actions are expressed in the same coordinate frame. The cumulative (or effective) symmetry at step $k$ is then defined as the closure of all these transported actions,
$\widehat{G}_{k}:=\big\langle\,G_{0}^{(k)},\,G_{1}^{(k)},\,\dots,\,G_{k-1}^{(k)}\,\big\rangle\subseteq\mathrm{Diff}(\Omega_{k}\times\Omega).$
Hence, each refinement augments the effective group by adjoining the new induced action $G_{k}^{(k)}$, yielding the extension $\widehat{G}_{k+1}=\langle\,\widehat{G}_{k},\,G_{k}^{(k+1)}\,\rangle$. This process captures how orthogonal, previously unmodelled symmetries become internalized during successive refinements.
The effective groups form an ascending chain of inclusions
$\widehat{G}_{0}\subseteq\widehat{G}_{1}\subseteq\widehat{G}_{2}\subseteq\dots,\quad\widehat{G}_{\infty}=\bigcup_{k\geq 0}\widehat{G}_{k}\subset\mathrm{Diff}(\Omega_{\infty}\times\Omega).$
Equivalently, the system $\{(\Omega_{k}\times\Omega,\,\widehat{G}_{k})\}_{k\geq 0}$ forms a direct (inductive) system of GG–spaces. Its colimit
[…]
represents the emergent symmetry acting on the limiting relation manifold. The sequence of intertwiners
$(\phi_{k}\times\mathrm{id}):(\Omega_{k}\times\Omega,\,\widehat{G}_{k})\longrightarrow(\Omega_{k+1}\times\Omega,\,\widehat{G}_{k+1})$
thus defines a coherent diagram of equivariant representations. The limiting group $\widehat{G}_{\infty}$ captures the symmetry closure induced by the hypothesized groups $\{G_{k}\}$—that is, the largest symmetry consistently representable within the chosen refinement sequence. When the input groups coincide with the true latent symmetries, $\widehat{G}_{\infty}$ recovers the full orbit structure of the data; otherwise it yields an approximate, model-dependent effective symmetry reflecting the inductive biases imposed during training.
In this way, the iterative refinement may be viewed as an algebraic growth process of symmetry representations, whose accuracy and expressiveness depend on the adequacy of the hypothesized groups.
The colimit view highlights how refinement accumulates symmetry structure. To make this process explicit, we next examine the algebraic relation between consecutive effective groups $\widehat{G}_{k}$ and $\widehat{G}_{k+1}$. Each step can be interpreted as an extension that measures how previously unmodelled, approximately orthogonal symmetries become integrated into the effective representation.
Group extensions and orthogonality.
At each iteration, the refinement augments the effective symmetry by adjoining the newly induced action $G_{k}^{(k)}$, yielding
$\widehat{G}_{k+1}=\langle\,\widehat{G}_{k},\,G_{k}^{(k)}\,\rangle\subseteq\mathrm{Diff}(\Omega_{k+1}\times\Omega).$
This refinement induces an extension of effective symmetry actions. When $\widehat{G}_{k}$ is normal in $\widehat{G}_{k+1}$, this extension can be represented by a short exact sequence
$1\rightarrow\widehat{G}_{k}\longrightarrow\widehat{G}_{k+1}\longrightarrow\Delta G_{k}\rightarrow 1,\qquad\Delta G_{k}:=\widehat{G}_{k+1}/\widehat{G}_{k},$
where $\Delta G_{k}$ measures the additional symmetry degrees of freedom introduced at refinement step $k$.
In the general case, $\Delta G_{k}$ should be interpreted as an effective quotient capturing new orbit directions modulo the previously internalized symmetries; formally, it may be viewed as a quotient in the category of group actions (equivalently, as a homogeneous space of $\widehat{G}_{k+1}$ modulo $\widehat{G}_{k}$), rather than as an abstract group.
When $\Delta G_{k}$ acts independently of $\widehat{G}_{k}$ (i.e. their intersection is trivial), the refinement is said to be orthogonal, and $\widehat{G}_{k+1}$ forms a semi–direct product $\widehat{G}_{k+1}\simeq\widehat{G}_{k}\rtimes\Delta G_{k}$. Otherwise, the groups overlap nontrivially, indicating that the newly added symmetry partially reuses or reparametrizes degrees of freedom already represented at earlier layers.
Hence, the refinement hierarchy can be understood as a sequence of nested symmetry extensions,
$1\to\widehat{G}_{0}\to\widehat{G}_{1}\to\widehat{G}_{2}\to\cdots,$
whose structure reflects how new symmetries are discovered and internalized during learning. Orthogonal extensions correspond to genuinely novel latent factors, whereas overlapping ones signal redundancy or coupling among the symmetry components already acquired.
In practice, orthogonal extensions correspond to the emergence of independent latent factors, each contributing a distinct and separable mode of variability within the representation. Non–orthogonal extensions, in contrast, indicate partial entanglement, where the newly introduced symmetry overlaps with or modulates those already internalized. From this perspective, iterative symmetry refinement can be viewed as a process of progressive factorization of the representation manifold, in which successive layers isolate increasingly independent symmetry components of the data–generating process. Whether this progressive separation succeeds depends on the adequacy of the hypothesized symmetry sequence $G_{k}$ and on the learnability of the associated intertwiners. When these conditions align, the refinement gradually decomposes composite transformations into independent symmetry factors, realizing disentanglement as an emergent structural property rather than a guarantee.
Regression via orbit completion
The previous subsection described the target-free case, where refinement is driven solely by hypothesized group symmetries. We now show how the same mechanism applies when an explicit mapping $f:\Omega\!\to\!Y$ is provided, yielding a geometric formulation of regression. When a target function $f:\Omega\to Y$ is given, its graph itself defines the relation manifold on which symmetry refinement operates.
Orbit completion and prediction.
Fix a target function $f:\Omega\to Y$ and its graph
$\Gamma_{f}=\{(x,f(x))\}\subset\Omega\times Y.$
Under refinement, we transport the graph by
$\begin{array}[]{rcl}\Gamma_{k}^{\prime}&=&\;\{\,(\Phi_{k}(x),\,f(x))\mid x\in\Omega\,\}\;\subset\;\Omega_{k+1}\times Y,\\[5.0pt] \Phi_{k}&=&\phi_{k}\circ\cdots\circ\phi_{0},\end{array}$
and hypothesize a group $G_{k}$ acting on $\Omega_{k+1}\times Y$ such that $\Gamma_{k}^{\prime}$ is (approximately) a single $G_{k}$–orbit. Let $a_{k}:=(\Phi_{k}(x_{\ast}),\,f(x_{\ast}))$ be an anchored training pair and write
$\mathcal{O}_{k}\;:=\;G_{k}\cdot a_{k}\;\subset\;\Omega_{k+1}\times Y.$
The refinement loss enforces $\Gamma_{k}^{\prime}\approx\mathcal{O}_{k}$ (orbit alignment). Given a novel input $x_{0}$, prediction amounts to completing the pair $(\Phi_{k}(x_{0}),\,y)$ so that it lies (as well as possible) on $\mathcal{O}_{k}$:
$\hat{y}_{k}(x_{0})\;\in\;\arg\min_{y\in Y}\;\mathcal{L}_{\mathcal{O},a_{k}}\bigr(\Phi_{k};G_{k},(\Phi_{k}(x_{0}),\,y)\bigr)$
Thus $y$ is the “missing puzzle piece” that places $(\Phi_{k}(x_{0}),y)$ on the current orbit.
Consistency across refinement.
As orbit alignment improves with $k$, $\mathcal{O}_{k}$ becomes an increasingly faithful proxy for $\Gamma_{k}^{\prime}$, so that $\hat{y}_{k}(x_{0})$ stabilizes. Under local transversality (the graph being a local section to the $G_{k}$–orbits) the completion is unique; otherwise, the minimum-norm or regularized solution is selected among admissible $y$’s.
Regression is realized implicitly: one learns $\Phi_{k}$ through symmetry refinement, and for a new input $x_{0}$, predicts $y$ by orbit completion. In this view, generalization emerges as the geometric continuity of orbit alignment across unseen samples.
Remark.
This perspective places supervised regression and unsupervised representation learning within the same algebraic geometry: both seek to embed data into a manifold of consistent orbits. Whereas the target-free case discovers stable orbits internally, the supervised case anchors them externally via the graph of $f$, revealing prediction as the act of completing a symmetry relation.
Theoretical Examples
Classes of actions on the latent space.
The general orbit–alignment principle subsumes classical invariance and equivariance as special cases of how the group acts on the target–latent product. If the group $G$ acts trivially on the target component, $g\!\cdot\!(x,y)=(gx,y)$, the orbit condition enforces $\phi(gx)=\phi(x)$ and the symmetry–loss reduces to the familiar invariant representation objective. If the target itself transforms under a representation $\rho:G\!\to\!\mathrm{Aut}(Y)$, i.e. $g\!\cdot\!(x,y)=(gx,\rho(g)y)$, orbit alignment becomes $\phi(gx)=\rho(g)\phi(x)$, recovering the classical notion of equivariance. More generally, when the transformation law couples input and output nontrivially—e.g. through a semi–direct or twisted product $1\!\to\!H\!\to\!G\!\to\!Q\!\to\!1$—the relation takes the form
$\phi(q\!\cdot\!x)=\rho(q)\phi(x)\tau(q,x),$
where the cocycle $\tau:Q\times\Omega\!\to\!H$ encodes the interaction between subgroups. Such “twisted” actions arise naturally when multiple latent factors interact hierarchically or multiplicatively. The refinement framework accommodates all these cases uniformly: each refinement step models a new extension $\widehat{G}_{k+1}=\langle\widehat{G}_{k},G_{k}^{(k+1)}\rangle$, thereby integrating additional coupling terms into the effective symmetry group without altering the underlying learning principle.
Remark.
The additional factor $\tau(q,x)\!\in\!H$ acts as a cocycle correction arising from the nontrivial extension $1\!\to\!H\!\to\!G\!\to\!Q\!\to\!1$. It encodes how transformations in the quotient $Q$ modulate those in the subgroup $H$. When $\tau\!\equiv\!e$, the law reduces to standard equivariance; a nontrivial $\tau$ corresponds to a twisted equivariance in which input transformations induce context–dependent modulations of the output.
Implementation and Computational Scalability
From a computational standpoint, each map $\phi_{k}$ is instantiated as a learnable neural intertwiner—typically a multilayer perceptron—trained to realize the corresponding group action through the proposed symmetry–loss. The key enabler of this framework is the formulation of the loss in terms of a complete fundamental system of symmetries $(\chi_{i})$, which provides an invariant coordinate system on the latent space. Optimization thus proceeds entirely in these invariant coordinates and does not require explicit evaluation or sampling of group transformations. This property makes the approach both scalable and adaptable: the group structure is captured implicitly through the algebraic dependencies among the $\chi_{i}$, rather than through explicit enumeration of its elements or representations. Consequently, the framework retains full architectural freedom—any differentiable model, such as a multilayer perceptron, convolutional network, or graph–based encoder, can serve as the intertwiner $\phi_{k}$. Training uses standard gradient–based optimizers (e.g. Adam or SGD) without modification, since the gradients are defined entirely through differentiable invariants. In this way, the refinement hierarchy becomes a neural ladder of intertwiners: a compositional stack of learned coordinate transformations that progressively internalize the symmetry structure of the data, combining theoretical expressiveness with computational efficiency.
Related Work and Conceptual Connections
The proposed framework builds upon and extends several converging lines of research on symmetry, representation learning, and neural computation.
This work can be viewed as a natural continuation of the initial orbit-alignment framework introduced in donmez2023discovering, which offered a first proof-of-concept but contained no refinement mechanism.
Early advances in disentangled representation learning emphasized the role of structured latent spaces that capture independent factors of variation (higgins2018towards; achille2018life). These approaches introduced the notion of latent homologies—correspondences between domains that share abstract relational structure—which closely parallels the present view of orbit alignment as matching transformations across representational levels. However, existing methods typically rely on explicit sampling or reconstruction-based objectives, whereas the present formulation leverages fundamental symmetries identified through invariant generators, leading to a symmetry–loss that enforces geometric consistency without requiring enumeration over group elements.
More recent theoretical work has explicitly connected symmetry-based representation learning with cognitive and neurobiological theories of general intelligence (Higgins2022). In this view, both artificial and biological systems achieve generalization by progressively internalizing the symmetries of their environment—a principle mirrored in the iterative refinement mechanism developed here. Our framework can thus be interpreted as a differentiable implementation of this symmetry-based learning process.
Within neuroscience, predictive-coding and free-energy principles provide complementary accounts of how the brain minimizes surprise by maintaining symmetry-consistent internal models (friston2009free; Cooray2024; Fumarola2022). The refinement steps in our model parallel hierarchical inference in cortical circuits, where successive layers align internal transformations with increasingly abstract invariances. This correspondence situates the proposed algorithm at the interface between geometric machine learning and computational neuroscience.
Finally, the developmental perspective articulated by Lake and colleagues (Lake2017)—that human learning proceeds through successive refinement of compositional and causal representations—offers a cognitive analogue to our iterative symmetry refinement: each $\phi_{k}$ learns to “re-align” internal coordinates with latent generative factors, akin to the child’s progressive discovery of invariant structure in perception and action. Together, these connections position the present framework as a unified geometric account of representation learning that spans artificial and biological domains.
Neurobiological Interpretation and Discussion
The iterative symmetry–refinement framework proposed here can be interpreted as a computational analogue of representational development in the brain. Biological learning does not begin with explicit supervision but through progressive internal reorganization of sensorimotor mappings to reflect the symmetries and regularities of the environment. From early visual processing to higher cognition, neural populations gradually acquire invariances—such as translation, reflection, or compositional structure— that stabilize perception and enable generalization (DiCarlo2012; Bertamini2018; Fumarola2022). This process can be viewed as a gradual alignment of internal symmetry groups to the dynamical symmetries of the external world (Higgins2022; Cooray2024).
Theoretical frameworks such as the Free–Energy Principle and Predictive Coding formalize this alignment as the minimization of surprise or prediction error (Friston2010; Parr2019). Within this view, our proposed refinement algorithm corresponds to a structural form of prediction minimization: rather than adjusting expectations about sensory states, it adapts the very symmetry relations that organize those expectations. Each refinement step thus parallels a neurodevelopmental process in which the brain learns to reorganize its internal representational geometry toward more abstract, factorized, and compositional forms.
This interpretation resonates with developmental theories of intuitive physics and learning–to–learn (Lake2017; marcus2020next). Infants appear to construct successively refined internal models that capture stable transformations in their sensorimotor world. In the same way that children infer latent causal variables underlying observed dynamics, each refinement step in our framework discovers latent symmetry generators that explain observed transformations. Analogously, each refinement step in our framework can be viewed as a learned “update of inductive biases,” internalizing new symmetry factors while preserving previously stabilized ones. In this sense, symmetry refinement provides a computational account of the gradual emergence of structured, symmetry–based intelligence— bridging neural development, representation learning, and theoretical neuroscience.
Outlook: Adaptive Symmetry Discovery
The present framework establishes symmetry refinement as a differentiable process that progressively internalizes structural regularities of the data. A natural next step is to couple this refinement mechanism with reinforcement learning principles, enabling the system to actively explore and hypothesize new symmetry transformations (caselles2019symmetry). Rather than fixing the candidate groups $\{G_{k}\}$ a priori, an agent could adjust its symmetry hypotheses based on predictive success or expected free–energy minimization, thereby learning which transformations yield stable orbits. This integration would extend the current model from passive symmetry alignment to an active symmetry discovery paradigm, in which perception, prediction, and action jointly shape the refinement trajectory. In such a view, representation learning, control, and curiosity become unified under the same organizing principle: the active pursuit of symmetry in the structure of experience.
From a neurobiological perspective, such a mechanism resonates with developmental learning in infants. Early sensorimotor exploration can be interpreted as a process of testing and refining internal symmetry hypotheses about the world— for instance, discovering invariances under translation, rotation, or causal reversal through embodied interaction (Lake2017; friston2019dynamic; Higgins2022). Each refinement step corresponds to an incremental reorganization of neural representations toward more abstract and compositional invariants, mirroring the progressive emergence of intuitive physics and structured world models in early cognition. Extending the proposed refinement algorithm with reinforcement signals would thus move it closer to the biological paradigm of learning to learn—where the discovery of symmetries is guided not only by reconstruction accuracy but also by the intrinsic drive to reduce uncertainty and maintain predictive control over the environment.
Acknowledgments
The author thanks his wife and children for their support and inspiration.